# Learning Network Representations with Disentangled Graph Auto-Encoder

### v1

@@ 4.3 Factor-wise Decoder
4.4 Evidence Lower Bound (ELBO)
Now we describe the objective of DVGA, following the paradigm of the VGAE (Kipf et al., 2016a). The model parameters, denoted as $\theta$, are acquired by maximizing the lower bound of log-likelihood, as specified in Eq.(11):
@@ 4.4 Evidence Lower Bound (ELBO)
$=$ $\mathop{\rm argmax}\limits_{\theta}\left[\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log\,p(A|Z)-D_{KL}(q_{\theta}(Z|X,A)||p(Z))\right]$ (24)
@@
$\log\,q_{\theta}(z_{u}|X,A)$ $=$ $\log\,q_{\theta}(z^{(0)}_{u}|X,A)-\sum_{m=1}^{M}\sum_{k=1}^{K}\log\,\left|\frac{{\rm d}f_{k}^{(m)}}{{\rm d}z_{u,k}^{(m-1)}}\right|$ (25)
$=$ $\log\,q_{\theta}(z^{(0)}_{u}|X,A)-\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{i=d^{\prime}+1}^{d}s_{i}(z_{u,k,1:d^{\prime}}^{(m-1)})$ (26)
@@
${\rm det}(J_{f^{(m)}}(z_{u,1}^{(m-1)},z_{u,2}^{(m-1)},\dots,z_{u,K}^{(m-1))}))$ $=$ $\left|\begin{array}[]{cccc}\frac{{\rm d}f_{1}^{(m)}}{{\rm d}z_{u,1}^{(m-1)}}&&&\\ &\frac{{\rm d}f_{2}^{(m)}}{{\rm d}z_{u,2}^{(m-1)}}&&\\ &&\ddots&\\ &&&\frac{{\rm d}f_{K}^{(m)}}{{\rm d}z_{u,K}^{(m-1)\mathstrut}}\end{array}\right|$ (32)
@@
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)-\log\,\left|{\rm det}\,\frac{\partial{z_{u}}}{\partial{z_{u}^{(0)}}}\right|$ (36)
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)-\sum_{m=1}^{M}\log{\rm det}\,\left|\frac{\partial{z_{u}^{(m)}}}{\partial{z_{u}^{(m-1)}}}\right|$ (37)
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)-\sum_{m=1}^{M}\sum_{k=1}^{K}\log\,\left|\frac{{\rm d}f_{k}^{(m)}}{{\rm d}z_{u,k}^{(m-1)}}\right|$ (38)
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)-\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{i=d^{\prime}+1}^{d}s_{i}(z_{u,k,1:d^{\prime}}^{(m-1)}).$ (39)
@@
In the above proof, the simplicity of the computations are attributed to the affine coupling flow used in Eqs. (20) and (21). In the end, after computing $\log\,q_{\theta}(z_{u}|X,A)$, we can use Eq. (3) and employ the reparameterization trick (Doersch, 2016) to get $D_{KL}(q_{\theta}(Z|X,A)||p(Z))$, so Eq. (24) can be calculated.
4.5 Statistical Independence of Mapping Subspaces
Strengthening the independence of the disentangled node representations explicitly contributes to the graph encoder’s ability to capture distinct information across various latent factors, potentially enhancing performance in downstream tasks. Next, we provide detailed explanations of independence regularization.
Remind our main goal is to empower the graph encoder to generate disentangled representations $z_{u}=\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$ for each latent factor. The $K$ different factors extracted by the dynamic routing mechanism are intended to focus on different connecting reasons, with each of the $K$ channels capturing mutually exclusive information related to latent factors. This emphasizes the importance of further promoting statistical independence among the disentangled representations to enhance the disentanglement. Therefore, before performing dynamic assignment mechanism, it is crucial to ensure that the focused perspectives of the $K$ subspaces in Eq. (14) are distinct. Hence, we consider imposing independent constraints on these $K$ subspaces.
Unfortunately, it is not trivial to obtain the solution that all disentangled subspaces are maximally different from each other. Therefore, we consider imposing independence constraints on the initial embedding vectors obtained from projections in $K$ different subspaces, as expressed in Eq. (14). Therefore, we approximate the solution by assigning unique labels to embedding vectors from the same subspace and optimizing the mapping subspaces as a vector classification problem. The discriminator, as represented in Eq. (40), is to distinguish which label a given embedding vector has:
@@ 4.5 Statistical Independence of Mapping Subspaces
This discriminator, denoted as $g_{\phi}$, comprises a fully connected neural network, and its parameters is $\phi$. It takes the embedding vector $c_{u,k}$ as inputs and generates a subspace-specific label $D_{k}$.
We design the following loss to train the discriminator:
@@
4.6 Optimization
Finally, our aim is to learn the parameters $\theta$ of the inference model and $\phi$ of the regularization term for DVGA within a unified framework. By employing gradient descent, we optimize the combined objective function, which integrates the Evidence Lower Bound (ELBO) and the independence regularizer:
Where $\lambda$ is a hyper-parameter that controls the balance between $-\mathcal{L}_{ELBO}(\theta)$ and the impact of the regularizer. $\mathcal{L}_{reg}(\phi)$ is the loss of the discriminator mentioned above, designed to encourage statistical independence among the disentangled node representations. The detailed training procedure for the DVGA model is illustrated in Algorithm 4.6.
@@ 4.6 Optimization
1:function DGCN($G=(V,E,X)$,$A$)
2:     for $u\in V$ do
3:         for $k=1,2,\dots,K$ do $\triangleright$ K channels.
4:              $c_{u,k}\leftarrow\frac{W_{k}^{\top}x_{u}+b_{k}}{\lVert W_{k}^{\top}x_{u}+b_{k}\rVert_{2}}$.
5:         end for
6:     end for
7:     for disentangle layer $l=1,2,\dots,L$ do
8:         $z_{u,k}^{t=1}\leftarrow c_{u,k},k=1,2,\dots,K,\forall u\in V$.
9:         for routing iteration $t=1,2,\dots,T$ do
10:              Calculate $p_{u,v}^{k,t-1},q_{v,u}^{k,t-1}$ by Eqs. (16) and (17).
11:              Update the $z_{u,k}^{t}$, $\forall u\in V$ by Eqs. (15) and (18).
12:         end for
13:         $c_{u,k}\leftarrow z_{u,k}^{t=T},k=1,2,\dots,K,\forall u\in V$$\triangleright$ when $l\leq L-1$.
14:     end for
15:     $z_{u}\leftarrow$ the concatenation of $z_{u,1},z_{u,2},\dots,z_{u,K}$.
16:     $Z\leftarrow$ each row represents the representation $z_{u}$ of the node $u$, $\forall u\in V$.
17:     return $Z$triangleright disentangled representation.
18:end function
19:function Encoder($G=(V,E,X)$,$A$)
20:     $\boldsymbol{\mu}$ = DGCN($G$,$A$),
21:     ${\rm log}\,\boldsymbol{\sigma}$ = DGCN($G$,$A$),
22:     $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\rm{I})$.
23:     $Z=\boldsymbol{\mu}+\boldsymbol{\epsilon}\circ\boldsymbol{\sigma}$. $\triangleright$ $Z$ is assigned the value of $\boldsymbol{\mu}$ at test time. The symbol $\circ$ denotes element-wise multiplication.
24:     return $Z$
25:end function
26:function Component-wise Flow($z_{u,k}$)
27:     for $m=1,2,\dots,M$ do
28:         Update $z_{u,k}$ by Eqs. (20) and (21).
29:     end for
30:     return $z_{u,k}$
31:end function
32:function Decoder($Z$)
33:     $Z_{k}\leftarrow$ the submatrix corresponding to the $k$-th channel in $Z$.
34:     $A_{factor}^{k}=\cos(Z_{k},Z_{k})\in\mathbb{R}^{N\times N}$.
35:     $\mathcal{D}(Z)\leftarrow maxpooling(A_{factor}^{1},A_{factor}^{2},\dots,A_{factor}^{K})$.
36:     $\hat{A}=\sigma(\mathcal{D}(Z)+ZZ^{\top})$.
37:     return $\hat{A}$
38:end function
39:for epoch $=1$ to $E$ do
40:     $Z,D_{KL}$ = Encoder($G$,$A$).
41:     $Z_{k}\leftarrow$ the submatrix corresponding to the $k$-th channel in $Z$.
42:     $Z_{k}$ = Component-wise Flow($\left\{z_{u,k}\right\}$), $Z_{k}=\left\{z_{u,k}\right\}$.
43:     $Z$= Stack($\left[Z_{1},Z_{2},\dots,Z_{K}\right]$, dim=0).
44:     $p(A|Z)=\textsc{Decoder}(Z)$.
45:     Calculate $D_{KL}(q_{\theta}(Z|X,A)||\mathcal{N}(\textbf{0},\textbf{I}))$ by Eq. (26).
46:     Calculate $\mathcal{L}_{ELBO}(\theta)$ by Eq. (11) and the independence regularizer $\mathcal{L}_{reg}(\phi)$ by Eq. (41).
47:     Update $\theta$ and $\phi$ to minimize $\mathcal{L}=-\mathcal{L}_{ELBO}(\theta)+\lambda\mathcal{L}_{reg}(\phi)$ by Eq. (42) .
48:end for
  For DGA, we employ the dynamic graph encoder in Section 4.1 to get latent representations, followed by the use of the factor-wise decoder in Eq. (22) for link prediction. The corresponding loss is given by:
@@
Where the first term corresponds to the reconstruction loss Eq. (13) for link prediction, typically implemented as a binary cross-entropy loss. The training process for the DGA model is outlined in Algorithm 4.6.
@@
5 Discussions
5.1 Time Complexity Analysis
@@
We have conducted a theoretical analysis of the time complexity of the proposed DVGA model, which is $O((K^{2}+fd+LT(C\frac{d}{K}+d))N+Md+(d+K)N^{2})=O((K^{2}+(f+LT)d+LTC)N+(d+K)N^{2}+Md)=O(N^{2})$. Here, $N$ denotes the number of nodes in the graph, $d$ is the dimensionality of latent variables, $f$ represents the dimensionality of node input features, $L$ is the number of layers in the disentangled layers employed by the dynamic disentangled graph encoder, $K$ indicates the number of latent factors and $T$ is the number of iterations for dynamic assignment. We define the maximum number of neighbors to be $C$. $d$, $T$, $L$, $C$ and $K$ are small constants. Specifically, the time complexity of our dynamic disentangled graph encoder is $O((fd+LT(c\frac{d}{K}+d))N)$. For the component-wise flow, $K$ channels are required to be implemented, each with a dimension of $\frac{d}{K}$, and $M$ steps, leading to a time complexity of $O(Md)$. The $O((d+K)N^{2})$ is attributed to the factor-wise decoder, indicating that the computational cost is predominantly affected by the size of the graph-structured data,as demonstrated in the empirical study presented in Section 6.6. Nevertheless, we argue that our proposed DVGA remains reasonably efficient in practical applications, particularly when considering the substantial performance improvements validated in our experiments. The time complexity of computing the regularizer in our model is $O(NK^{2})$. Similarly, we are able to get that the time complexity of DGA is $O((K^{2}+fd+LT(C\frac{d}{K}+d))N+(d+K)N^{2})=O((K^{2}+(f+LT)d+LTC)N+(d+K)N^{2})=O(N^{2})$.
5.2 Number of Parameters Analysis
@@
The number of parameters for the proposed DVGA is $O(fd+KMd^{\prime}(d-d^{\prime}))=O(fd+KMd^{2})$, where $N$ is the number of nodes, $f$ represents the dimensionality of node input features, $K$ is the number of latent factors, $d$ is the dimensionality of latent variables and $d^{\prime}$ is employed in the affine coupling parameters of flows to determine the dimensions that remain unchanged from the previous step to the subsequent one. Specifically, the number of parameters for the dynamic disentangled encoder is $O(fd)$, the flow utilizes $O(KMd^{\prime}(d-d^{\prime}))$ parameters, and the independence regularization term involves $O(d)$ parameters. This parameter complexity is independent of the number of nodes and edges in the input graph, resulting in a relatively small value. Likewise, the number of parameters in DGA without utilizing flow is $O(fd)$.
6 Experiments
We show the performance of our proposed models and demonstrate their effectiveness in this section by presenting results on two unsupervised graph analytic tasks, including link prediction and node clustering. Additionally, our models are evaluated on a semi-supervised graph analytic task, namely, node classification. We also qualitatively investigate our models’ performance in factor disentanglement. Further, we analyze the convergence behavior of our models and provide module ablation studies along with parameter sensitivity analysis.
6.1 Link prediction
@@
Link prediction aims to predict whether there will be an edge or not between a pair of nodes (Loehlin, 2004). It stands as one of the most prevalent tasks in representation learning of network. Utilizing the decoder of our model, we can effectively recover lost links or predict potential links through an analysis of the given graph.
@@
(1) Real-World Datasets: We assess our models performance on three widely used benchmark citation network datasets: Cora, CiteSeer, and PubMed, where nodes correspond to research papers, edges represent citations, and labels indicate the research areas (Sen et al., 2008). The features at the node level correspond to textual attributes found in the papers. Additionally, each node is assigned with a label indicating its class. We treat the graphs as undirected graphs. Table 1 provides a summary of the dataset statistics.
Table 1: Statistics for the citation network datasets.
@@
(2) Synthetic Dataset: To explore the performance of our models on graphs with $m$ latent factors, we follow the methodology from Guo et al. (2022) for constructing synthetic graphs. Specifically, we create $m$ Erdös-Rényi random graphs, each comprising 1000 nodes and 16 classes. Nodes are connected with a probability $p$ within the same class, and nodes are connected with a probability $q$ within different classes. Subsequently, by summing their adjacency matrices, we aggregate these generated graphs and set elements greater than zero to one. Then , we get the final synthetic graph with $m$ latent factors. This composite graph includes $16m$ classes, and assigns $m$ labels to each node, derived from the original labels of $m$ random graphs. Node representations are established based on the rows of the adjacency matrix. We fix the value of $q$ at $3e^{-5}$ and fine-tune $p$ to attain an average neighborhood size of around 40.
@@
(3) Baselines: We evaluate the performance of DVGA and DGA in comparison to state-of-the-art baselines, including both graph embedding methods, such as Spectral Clustering (SC) (Tang et al., 2011), DeepWalk (Perozzi et al., 2014), and node2vec (Grover et al., 2016), as well as graph auto-encoder algorithms, including GAE/VGAE (Kipf et al., 2016a), ARGA/ARVGA (Pan et al., 2018), LGAE/LVGAE (Salha et al., 2021), SIGVAE (Hasanzadeh et al., 2019) and GNAE/VGNAE (Ahn et al., 2021). Notably, SC, DeepWalk and node2vec lack the capability to use node features during the embedding learning process. Therefore, we evaluate them only on datasets without node features.
@@
(4) Evaluation metrics: We employ two metrics, Area Under the ROC Curve (AUC) and Average Precision (AP), to evaluate performance (Kipf et al., 2016a). Each experiment is conducted five times, and the final scores are reported as mean values with standard errors. The real-world datasets are partitioned into training, testing, and validation sets. We augment the original graph with a balanced set of positive and negative (false) edges in the validation and testing sets. Specifically, the validation set comprises 5% edges, the test set includes 10% edges, and the remaining edges are used for training. Both the validation and testing sets include an equal number of non-edges. For synthetic datasets, we perform a random split, allocating 60% to the training set, 20% to the validation set, and another 20% to the test set (Guo et al., 2022).
@@
(5) Parameter Settings: We configure our DGA and DVGA models with $d=K\Delta d$, where $d$ represents the hidden dimension, $K$ denotes the number of channels, and $\Delta d=\frac{d}{K}$ is the output dimension of each channel. Hyper-parameter tuning is performed on the validation split of each dataset using Optuna (Akiba et al., 2019) for efficiency. Specifically, Optuna is executed for 100 trials for each configuration, and the hyper-parameter search space is defined as follows: dropout $\sim\{0,0.05,...,1\}$ with a step of 0.05, learning rate $\sim[1e^{-3},1]$, the number of channels $K\in\{2,3,...,10\}$, the iteration of routing $T\in\{1,2,...,10\}$, the number of layers $L\in\{1,2,...,6\}$, and the regularization term $\lambda\sim[1e^{-5},1]$. The output dimension of each channel is $\Delta d=16$. For DVGA, the number of flow steps $\in\{1,2,...,5\}$. Using the best hyper-parameters, the models are trained for 3000 epochs, and the performance is reported based on five runs on the test split. Regarding the remaining baselines, we maintain consistency with our models in terms of the number of latent dimensions, while other parameters follow the configurations described in the corresponding papers.

### v2

@@ IV-C Factor-wise Decoder
IV-D Evidence Lower Bound (ELBO)
Now we describe the objective of DVGA. The model parameters, denoted as $\theta$, are obtained by maximizing the lower bound of log-likelihood, as specified in Eq.(10):
@@ IV-D Evidence Lower Bound (ELBO)
$=$ $\mathop{\rm argmax}\limits_{\theta}\left[\mathbb{E}_{Z\sim q_{\theta}(Z|X,A)}\log\,p(A|Z)\right.$
$\left.-D_{KL}(q_{\theta}(Z|X,A)||p(Z))\right]$ (23)
@@
$\log\,q_{\theta}(z_{u}|X,A)=$ $\log\,q_{\theta}(z^{(0)}_{u}|X,A)$
$-\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{i=d^{\prime}+1}^{d}s_{i}(z_{u,k,1:d^{\prime}}^{(m-1)})$ (24)
@@
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)$
$-\sum_{m=1}^{M}\sum_{k=1}^{K}\log\,\left|\frac{{\rm d}f_{k}^{(m)}}{{\rm d}z_{u,k}^{(m-1)}}\right|$ (29)
$=$ $\log\,q_{\theta}(z_{u}^{(0)}|X,A)$
$-\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{i=d^{\prime}+1}^{d}s_{i}(z_{u,k,1:d^{\prime}}^{(m-1)}).$ (30)
@@
In the above proof, the simplicity of the computations are attributed to the affine coupling flow used in Eqs. (19) and (20). In the end, after computing $\log\,q_{\theta}(z_{u}|X,A)$, we can use Eq. (3) and employ the reparameterization trick [58] to get $D_{KL}(q_{\theta}(Z|X,A)||p(Z))$, so Eq. (23) can be calculated.
IV-E Statistical Independence of Mapping Subspaces
Our main goal is to empower the graph encoder to generate disentangled representations $z_{u}=\left[z_{u,1},z_{u,2},\dots,z_{u,K}\right]$ for each latent factor. The $K$ different factors extracted by the dynamic routing mechanism are intended to focus on different connecting causes, with each of the $K$ channels capturing mutually exclusive information related to every latent factors. This highlights the significance of encouraging statistical independence among the disentangled representations in order to improve the process of disentanglement. Therefore, it is crucial to ensure that the focused perspectives of the $K$ subspaces in Eq. (13) are distinct before performing dynamic assignment mechanism. Hence, we propose imposing independent constraints on these $K$ subspaces.
Unfortunately, it is not trivial to obtain the solution that all disentangled subspaces are maximally different from each other. Therefore, we propose imposing independence constraints on the initial embedding vectors obtained from projections in $K$ different subspaces, as expressed in Eq. (13). We estimate the solution by giving distinct labels to embedding vectors belonging to the same subspace and maximizing the mapping subspaces as a problem of vector classification. The discriminator, as denoted by the Eq. (31), is used to differentiate the label associated with a particular embedding vector:
@@ IV-E Statistical Independence of Mapping Subspaces
This discriminator, represented as $g_{\phi}$, comprises a fully connected neural network, and its parameters are denoted as $\phi$. It takes the embedding vector $c_{u,k}$ as inputs and generates a subspace-specific label $D_{k}$.
We formulate the following loss function to train the discriminator:
@@
Explicitly strengthening the independence of the disentangled node representations will improve the graph encoder’s ability to capture different information across several latent factors, potentially boosting performance in subsequent tasks.
@@
1:function Encoder($G=(V,E,X)$,$A$)
2:     $\boldsymbol{\mu}$ = DGCN($G$,$A$),
3:     ${\rm log}\,\boldsymbol{\sigma}$ = DGCN($G$,$A$),
4:     $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\rm{I})$.
5:     $Z=\boldsymbol{\mu}+\boldsymbol{\epsilon}\circ\boldsymbol{\sigma}$. $\triangleright$ $Z$ is assigned the value of $\boldsymbol{\mu}$ at test time. The symbol $\circ$ denotes element-wise multiplication.
6:     return $Z$
7:end function
8:function Component-wise Flow($z_{u,k}$)
9:     for $m=1,2,\dots,M$ do
10:         Update $z_{u,k}$ by Eqs. (19) and (20).
11:     end for
12:     return $z_{u,k}$
13:end function
14:function Decoder($Z$)
15:     $Z_{k}\leftarrow$ the submatrix corresponding to the $k$-th channel in $Z$.
16:     $A_{factor}^{k}=\cos(Z_{k},Z_{k})\in\mathbb{R}^{N\times N}$.
17:     $\mathcal{D}(Z)\leftarrow maxpooling(A_{factor}^{1},A_{factor}^{2},\dots,A_{factor}^{K})$.
18:     $\hat{A}=\sigma(\mathcal{D}(Z)+ZZ^{\top})$.
19:     return $\hat{A}$
20:end function
21:for epoch $=1$ to $E$ do
22:     $Z,D_{KL}$ = Encoder($G$,$A$).
23:     $Z_{k}\leftarrow$ the submatrix corresponding to the $k$-th channel in $Z$.
24:     $Z_{k}$ = Component-wise Flow($\left\{z_{u,k}\right\}$), $Z_{k}=\left\{z_{u,k}\right\}$.
25:     $Z$= Stack($\left[Z_{1},Z_{2},\dots,Z_{K}\right]$, dim=0).
26:     $p(A|Z)=\textsc{Decoder}(Z)$.
27:     Calculate $D_{KL}(q_{\theta}(Z|X,A)||\mathcal{N}(\textbf{0},\textbf{I}))$ by Eq. (24).
28:     Calculate $\mathcal{L}_{ELBO}(\theta)$ by Eq. (10) and the independence regularizer $\mathcal{L}_{reg}(\phi)$ by Eq. (32).
29:     Update $\theta$ and $\phi$ to minimize $\mathcal{L}=-\mathcal{L}_{ELBO}(\theta)+\lambda\mathcal{L}_{reg}(\phi)$ by Eq. (33) .
30:end for
IV-F Optimization
Finally, our aim is to learn the parameters $\theta$ of the inference model and $\phi$ of the regularization term for DVGA in a unified framework. By employing gradient descent, we optimize the combined objective function that integrates the Evidence Lower Bound (ELBO) and the independence regularizer:
where $\lambda$ is a hyper-parameter that controls the balance between $-\mathcal{L}_{ELBO}(\theta)$ and the impact of the regularizer. $\mathcal{L}_{reg}(\phi)$ is the loss of the discriminator, designed to encourage statistical independence among the disentangled node representations. The detailed training procedure for the DVGA model is shown in Algorithm 1.
For DGA, we employ the dynamic graph encoder in Section IV-A to get latent representations, followed by the use of the factor-wise decoder in Eq. (21) for link prediction. The corresponding loss is given by:
@@ IV-F Optimization
where the first term corresponds to the reconstruction loss Eq. (12) for link prediction, typically implemented as a binary cross-entropy loss. The training process for the DGA model is outlined in Algorithm 2.
@@
V Discussions
V-A Time Complexity Analysis
@@
We have conducted a theoretical analysis of the time complexity of the proposed DVGA model, which is $O((K^{2}+fd+LT(C\frac{d}{K}+d))N+Md+(d+K)N^{2})=O((K^{2}+(f+LT)d+LTC)N+(d+K)N^{2}+Md)=O(N^{2})$. Here, $N$ denotes the number of nodes in the graph, $d$ is the dimensionality of latent variables, $f$ represents the dimensionality of node input features, $L$ is the number of layers in the disentangled layers employed by the dynamic disentangled graph encoder, $K$ indicates the number of latent factors and $T$ is the number of iterations for dynamic assignment. We define the maximum number of neighbors to be $C$. $d$, $T$, $L$, $C$ and $K$ are small constants.
Specifically, the time complexity of our dynamic disentangled graph encoder is $O((fd+LT(c\frac{d}{K}+d))N)$. For the component-wise flow, $K$ channels are required to be implemented, each with a dimension of $\frac{d}{K}$, and $M$ steps, leading to a time complexity of $O(Md)$. The $O((d+K)N^{2})$ is attributed to the factor-wise decoder, indicating that the computational cost is predominantly affected by the size of the graph-structured data, as demonstrated in the empirical study presented in Section VI-F. Nevertheless, we argue that our proposed DVGA remains reasonably efficient in practical applications, particularly when considering the substantial performance improvements validated in our experiments. The time complexity of computing the regularizer in our model is $O(NK^{2})$.
Similarly, we are able to get that the time complexity of DGA is $O((K^{2}+fd+LT(C\frac{d}{K}+d))N+(d+K)N^{2})=O((K^{2}+(f+LT)d+LTC)N+(d+K)N^{2})=O(N^{2})$.
V-B Number of Parameters Analysis
The number of parameters for the proposed DVGA is $O(fd+KMd^{\prime}(d-d^{\prime}))=O(fd+KMd^{2})$, where $d^{\prime}$ is employed in the affine coupling parameters of flows to determine the dimensions that remain unchanged from the previous step to the subsequent one.
@@ V-B Number of Parameters Analysis
Specifically, the number of parameters for the dynamic disentangled encoder is $O(fd)$, the flow utilizes $O(KMd^{\prime}(d-d^{\prime}))$ parameters, and the independence regularization term involves $O(d)$ parameters. This parameter complexity is independent of the number of nodes and edges in the input graph, resulting in a relatively small value.
Likewise, the number of parameters in DGA is $O(fd)$.
VI Experiments
In this section, we compare our method with other state-of-the-art methods in three different applications, including unsupervised learning for link prediction and node clustering, as well as semi-supervised learning for node classification. We also qualitatively investigate our models’ performance in factor disentanglement.
VI-A Link prediction
@@
Link prediction aims to predict the presence or absence of a connection between two nodes [1]. It is one of the most common tasks in the representation learning of networks. By using the decoder in our model, we can effectively recover missing links or predict potential links through an analysis of the given graph.
@@
1) Real-World Datasets: We evaluate the effectiveness of our models on three widely used benchmark citation network datasets: Cora, CiteSeer, and PubMed. In these datasets, nodes correspond to research papers, edges represent citations, and labels indicate the research areas [63]. The features at the node level correspond to textual attributes found in the papers. Additionally, each node is assigned with a label indicating its class. We treat the graphs as undirected graphs. Table I provides a summary of the dataset statistics.
TABLE I: Statistics for the citation network datasets.
@@
2) Synthetic Dataset: To explore the performance of our models on graphs with $m$ latent factors, we follow the methodology from [46] for constructing synthetic graphs. Specifically, we create $m$ Erdös-Rényi random graphs, each consisting of 1000 nodes and 16 classes. Nodes within the same class are connected with a probability of $p$, while nodes from other classes are connected with a probability of $q$. Subsequently, by summing their adjacency matrices, we aggregate these generated graphs and set elements greater than zero to one. Then, we get the final synthetic graph with $m$ latent factors. This composite graph includes $16m$ classes and assigns $m$ labels to each node, which are obtained from the original labels of $m$ random graphs. Node representations are established based on the rows of the adjacency matrix. We fix the value of $q$ at $3e^{-5}$ and fine-tune $p$ to attain an average neighborhood size of around 40.
@@
3) Baselines: We evaluate the performance of DVGA and DGA in comparison to state-of-the-art baselines, including graph embedding methods, such as Spectral Clustering (SC) [64], DeepWalk [24], and node2vec [26], as well as graph auto-encoder algorithms, including GAE/VGAE [12], ARGA/ARVGA [23], LGAE/LVGAE [37], SIGVAE [14] and GNAE/VGNAE [65]. Notably, SC, DeepWalk and node2vec lack the capability to use node features during the embedding learning process. Therefore, we evaluate them only on datasets without node features.
@@
4) Evaluation metrics: We employ two metrics, Area Under the ROC Curve (AUC) and Average Precision (AP), to evaluate the models’ performance [12]. Each experiment is conducted five times, and the final scores are reported as mean values with standard errors. The real-world datasets are partitioned into training, testing, and validation sets. We augment the original graph with a balanced set of positive and negative (false) edges in the validation and testing sets. Specifically, the validation set comprises 5% edges, the test set includes 10% edges, and the remaining edges are used for training. Both the validation and testing sets include an equal number of non-edges. For synthetic datasets, we perform a random split, allocating 60% to the training set, 20% to the validation set, and another 20% to the test set [46].
@@
5) Parameter Settings: We configure our DGA and DVGA models with $d=K\Delta d$, where $d$ represents the hidden dimension, $K$ denotes the number of channels, and $\Delta d=\frac{d}{K}$ is the output dimension of each channel. Hyper-parameter tuning is performed on the validation split of each dataset using Optuna [66] for efficiency. Specifically, Optuna is executed for 100 trials for each configuration, and the hyper-parameter search space is defined as follows: dropout $\sim\{0,0.05,...,1\}$ with a step of 0.05, learning rate $\sim[1e^{-3},1]$, the number of channels $K\in\{2,3,...,10\}$, the iteration of routing $T\in\{1,2,...,10\}$, the number of layers $L\in\{1,2,...,6\}$, and the regularization term $\lambda\sim[1e^{-5},1]$. The output dimension of each channel is $\Delta d=16$. For DVGA, the number of flow steps $\in\{1,2,...,5\}$. The models are trained for 3000 epochs using the optimal hyper-parameters, and the performance is evaluated based on five test split runs. For the remaining baselines, we maintain consistency with our models in terms of the number of latent dimensions, while other parameters are set according to the configurations stated in the corresponding papers.